OpenStudy (anonymous):

Make a t-table of values for y = x2 + 4x + 1. A suggestion is to begin your values with -5. 1.List your ordered pairs from the table. Use parenthesis for each pair, such as (-1,2). 2.Which point is the vertex of the parabola? a) (-2, -3) b) (2, 3) c) (-2, 3) d) (2, -3) 3. The line which you folded is called the line of symmetry. This divides the parabola into two matching halves or mirror images. The equation of the line is x = ___ because it is vertical. Look to see where your folded line crosses the x-axis. Choose the equation below. a) x = -2 b) x = 2 c) x = 0 d)None

OpenStudy (anonymous):

You can write a table on your own. On to the vertex; When an equation is written as: \[f(x)=ax^2+bx+c\] The vertex of a parabola will be the point: \[(-\frac{b}{2a},f(-\frac{b}{2a}))\] \[Vertex (-\frac{4}{2(1)},f(-\frac{4}{2(1)}))=(-2, f(-2))=(-2,-3)\] To find the line of symmetry, you can rewrite the equation to this form and the line of symmetry will be at x=h: \[f(x)=a(x-h)^2+K\] To do this, you need to complete the square for x. \[ y = x^2 + 4x + 1\] \[y=1(x^2+4x+4-4+1)=(x+2)^2-3\] From this we can see that h=-2. The line of symmetry is at x=h=-2.